### v1

# Quantum determinism and completeness restored by indistinguishability and long-time particle detection

Abstract
We argue that measurement data in quantum physics can be rigorously interpreted only as a result of a statistical, macroscopic process, taking into account the indistinguishable character of identical particles. Quantum determinism is in principle possible on the condition that a fully-fledged quantum model is used to describe the measurement device in interaction with the studied object as one system. In contrast, any approach that relies on Born’s rule discriminates the dynamics of a quantum system from that of the detector with which it interacts during measurement. In this work, we critically analyze the validity of this measurement postulate applied to single-event signals. In fact, the concept of “individual” particle becomes inadequate once both indistinguishability and a scattering approach allowing an unlimited interaction time for an effective detection, are considered as they should be, hence preventing the separability of two successive measurement events. In this context, measurement data should therefore be understood only as a result of statistics over many events. Accounting for the intrinsic noise of the sources and the detectors, we also show with the illustrative cases of the Schrödinger cat and the Bell experiment that once the Born rule is abandoned on the level of a single particle, realism, locality and causality are restored. We conclude that indiscernibility and long-time detection process make quantum physics not fundamentally probabilistic.
I Introduction
Quantum mechanics is in its original formulation compatible with any experimental result obtained using a statistics-based approach for the measurement of a quantum system’s properties. However, the information that can be extracted from measurement data being fundamentally limited and prone to interpretation biases, complete information on a system’s quantum state remains inaccessible to human knowledge. Besides complementarity and the Heisenberg uncertainty relations, which impose absolute bounds on the precision of the simultaneous measurement of pairs of observables, the main limiting factor is that no quantum state is immune to a measurement process, and that the obtained data rather reflects the state of the system in interaction with the measurement device. This fact has been extensively discussed in relation with the no-cloning theorem [1, 2].
To overcome the difficulty of giving an interpretation of the measurement data at the level of elementary particles without an exact, detailed knowledge of the experimental conditions on the source and the detector (thus involving a macroscopic number of particles), full determinism, which is seen as a cumbersome obstacle, has been abandoned in favor of probabilistic predictability with a set of postulates, including the Born rule [3]. One major consequence is that in this framework, properties of quantum systems are assumed to acquire a definite value only because of measurement, thereby implying that there cannot be an ontological continuity between classical and quantum physics: quantum properties become physical quantities only when they are measured. The uncertainty principle makes the matter even more intricate as the precise measurement of one quantum variable (e.g., momentum) renders impossible for its conjugate quantity (position) the acquisition of a definite value. But since the postulates provide a formal ground (based on axioms formulated by Dirac[4] and von Neumann [5]) for both the description of a quantum system and the measurement of its properties in terms of states, observables and their time-evolution, quantum mechanics is commonly taken as an operatorial pragmatic approach based on a set of postulates for measurement and control. An essential question remains though: whether physics at the quantum level fundamentally is probabilistic or deterministic.
Here, we argue that the Born rule (or probability measurement postulate) is a non-essential mathematical convenience that should no longer be used, even though for nearly 100 years, much research in quantum physics has relied upon it for the interpretation of quantum experimental data. In a nutshell, this rule states that the probability density of finding a system in a given state, immediately after measurement, is proportional to the square of the amplitude of the system’s wave function in that state immediately before measurement, thus assuming that the system was isolated from its environment during the measurement process. In this framework, the measurement of a physical quantity can only yield one of the eigenvalues of the corresponding observable with a certain probability. As such, the Born rule separates the dynamics of the studied quantum system from that of the measurement device, thus assuming that measurement data correspond to observables as properties of the studied quantum system only. However, the elephant in the room that has been systematically overlooked in practice, is that there is no reason why quantum mechanics should discriminate between the object of study and the measurement device it is in interaction with as though they were independent entities. A precise knowledge of the interaction between the detector and the quantum system is required for a detailed and rigorous description of the measurement process and hence a sound interpretation of the data. This is a considerable challenge. Another difficulty is the implicit use of the concept of probability that does not appear anywhere during the quantum dynamics but only at the measurement instance, the exact time and the duration of which cannot be rigorously known.
The application of the Born rule has become prevalent for the probabilistic interpretation of measurement outcomes. While this facilitates the use of classical concepts and language to describe quantum phenomena and properties, the trade-off is in the necessity to consider the measurement device as classical. Ceasing to apply the Born rule would allow to discard the notion of probabilities being a constitutive element of quantum mechanics thereby paving the way to its deterministic interpretation. The cost would then be a more elaborate description of the detection process. Given the widespread use of the theory in its probabilistic interpretation and its achievements, one may argue that while not minor, the matter remains a specialized problem with no strong consequence for experiments as discussed in a recent editorial [6]. However, Werner Heisenberg’s assertion that: “What we observe is not nature in itself but nature exposed to our method of questioning[7], can serve as a warning against the systematic use of any methodology that eludes fundamental difficulties to interpret experimental data.
The present work is devoted to the analysis of measurement in quantum physics experiments to provide a more adequate view of the theory by avoiding its probabilistic interpretation. The article is organized as follows. In Section II, we critically discuss the Born rule as a corner stone of most of the research in quantum physics. We show that not only it is possible to dismiss it but that it must be abandoned for the modeling of single-event measurements. In Section III, we discuss key aspects of quantum physics experiments, notably that they are fundamentally based on statistical processes. To substantiate our critical discussion, we provide two illustrative examples: The Schrödinger cat experiment and the Bell experiment in Sections IV and V respectively. The article ends with a discussion and concluding remarks.
II The inadequacy of the Born rule for single events
At the time of the formulation of the Born rule in 1926 [3], physicists and chemists ambitioned and managed to gain a detailed knowledge and understanding of the electronic structures of atoms and molecules, neglecting the effects of the apparatus upon measurement of a system’s quantum properties. Contributions of other authors, including von Neumann[5], Dirac[8], and Jordan[9, 10], which adhere to the probabilistic interpretation of quantum mechanics, made the Born rule a postulate of quantum mechanics [11]. As this rule is still used nowadays, the dichotomy between the object of study and the measurement device persists; and it is legitimate to a why a detector during measurement is not described as a quantum system together with the object of study. Moreover, for single microscopic event (or individual) measurements, the use of this rule is questionable.
II.1 Redefinition of the measurement without the Born rule
Any measurement is in essence a macroscopic process and the use of the Born rule is unnecessary as long as a suitable description of the detection process is done. Generally, two key aspects of measurement are neglected when performing an experiment and analyzing the data.
First, the identification and spatial isolation of an elementary particle or an atom is impossible to achieve since being either a boson or a fermion, it is indistinguishable from other identical particles or atoms in an experiment, especially when its energy is below that which is required for an excitation process in a detector to be effective. Typically, the threshold is of the order of 1 eV to excite an electron in a semiconductor-based detector. Note that the so-called indistinguishability of identical particles (including composite ones like atoms or ions) is one way to consider that ultimately only quanta or excitations of quantum fields exist with specific quantum numbers in contrast to point-like particles that could be tagged individually.
To illustrate this important point, we consider two localized states and rewrite them as a delocalized state. For example, suppose two local points $1$ and $2$ where we define the cat state: $|\Psi\rangle=[(\hat{a}^{\dagger}_{1})^{2}-(\hat{a}^{\dagger}_{2})^{2}]|0\rangle/2$ where $\hat{a}^{\dagger}_{i}$ is a creation operator at point $i=1,2$ acting on the vacuum state $|0\rangle$. We immediately see that the rewriting: $|\Psi\rangle=({\hat{a}^{\dagger}_{1}}+{\hat{a}^{\dagger}_{2}})({\hat{a}^{\dagger}_{1}}-{\hat{a}^{\dagger}_{2}})|0\rangle/2$ can be interpreted as a factorized combination of two orthogonal symmetric and anti-symmetric delocalized single photon states. Conversely, the state $|\Psi\rangle={\hat{a}^{\dagger}_{1}}\hat{a}^{\dagger}_{2}|0\rangle=[({\hat{a}^{\dagger}_{1}}+{\hat{a}^{\dagger}_{2}})^{2}-({\hat{a}^{\dagger}_{1}}-{\hat{a}^{\dagger}_{2}})^{2}]|0\rangle/4$ can be seen as two localized separated states or a cat state of two delocalized states. This observation is a consequence of indiscernability and enforces the use of statistics to describe a quantum state as a superposition of different elementary particle numbers. Even in the absence of an interaction force, bosons (fermions) in beams attract (repel) each other resulting in (anti)bunching effect as a result of their mutual quantum correlations. Therefore, an essential improvement in a theoretical description is to use a multi-particle and multi-mode state in any scattering process rather than the usual amplitude probability associated with a state containing a determined particle number. Even if in an experiment bosons would be separated far apart from each other in an anti-bunching (or classical) regime, their indistinguishable nature remains and guarantees the impossibility of separating a quantum state into a single particle wave-function in a unique manner as established from the quantum field formalism [12, 13].
Second, in practice one hardly ponders over the significance of measuring an observable whose prediction by quantum mechanics is such that its mean-square deviation provides an error bar comparable to its average (expected value). Normally, one measures quantities with a relatively small confidence interval, otherwise one cannot assess the reliability of a measurement process outcome. Since any EPR or Bell states experiment leads to noisy outcomes, the feasibility of a detection that discriminates between the absence or the presence of a particle raises questions (which will be discussed further below). An avalanche diode works realistically once the uncertainty of having more than one photon interacting is lower than the uncertainty associated with the photon coming from the ambient thermal noise. So, we need the window time $T$ to be large enough to reduce the uncertainty associated with the photon number entering into the detector. For these reasons, we propose that approaches based on the Born rule be replaced by another more pragmatic rule:
A measurement is reliable if its outcome has a quantum uncertainty that is much smaller than the uncertainty associated to its intrinsic noise, mostly the thermal detector noise.
As a result of the two observations discussed above, there is no need to determine a probability amplitude anymore but instead any macroscopic quantity related to the multi-particle state. Since statistics are necessary but can be performed without use of any probability or stochastic approach, one may suggest alternatively that any experiment is described in terms of macroscopic quantities as done in thermodynamics. A macroscopic quantity such as an accumulated charge due to an electric current, falls in this category. As a consequence, an accurate detector description would include the quantum process of any signal amplification to a macroscopic quantity involving a large ensemble of particles whose relative fluctuations shrink to zero in the thermodynamic limit. A macroscopic approach implies the use of many quantum modes involving a superposition of Fock states. This statement is also valid for any higher-order moment of any macroscopic observable like the fluctuations of particle number or macroscopic current. On the contrary, considering one mode, the particle number and its energy are fluctuating. A particle mode interacting with a detector spreads fluctuations inside the detector rendering a scattering less effective and more difficult to model and to interpret. In the EPR experiment, we show further below that it is the difference between two outcomes with negligible fluctuations that appears to be measurable, thus rendering the quantum measurement local again.
II.2 Deterministic view of quantum mechanics
The renouncement to a probabilistic interpretation as an intrinsic feature of quantum mechanics leads to reassess more firmly the deterministic view of quantum mechanics albeit with a different interpretation from the one given in the EPR experiment [14, 15, 16, 17, 18]. But in contrast to classical mechanics, in a deterministic view, any quantum observable like position, momentum or particle number is not well defined for microscopic systems, and so far only the quantum wave function makes sense. Note though that they can be used in a macroscopic or coarse grained sense when the relative uncertainty can be made sufficiently small. In this sense, a precision in the position of $10^{-2}$Å (as in an electronic microscope experiment) is possible for a macroscopic object but not for an individual atom, unless it is chemically bonded to the former (though while still individual as an impurity, it is no longer isolated). Probability distributions at the microscopic level should therefore be interpreted only in the statistical sense, i.e. viewed as a histogram over many events.
In this deterministic paradigm, the principle of realism amounts to saying that the measured macroscopic quantity has vanishing fluctuations. It sounds a priori idealistic that a perfect exact measurement requires an infinite number of interrogations but, by analogy, it sounds as idealistic to assume that space and time are continuous since one never approaches an infinitesimal distance or time increment. One needs only to accept that, in any modeling, concepts are defined only in some unreachable limit cases; otherwise, this would add another layer of unwanted complexity preventing a simple understanding of a physical phenomena. This corroborates the non-cloning theorem proposed by Zurek [1] on the impossibility to clone any unknown quantum state justifying the use of many copies to gain knowledge of this state before reproducing it. We therefore propose that: Quantum mechanics predicts deterministically that the humankind cannot predict everything.
By abandoning the Born rule, the quantum theory becomes complete with no quantum randomness. In fact, incompleteness is rooted in the addition of Born’s rule as a postulate, which imposes itself in quantum mechanics as relevant and useful to interpret experimental data using the unjustified simplification that a macroscopic experiment is the result of many separable individual experiments. If this assumption seems to work in practice, it neglects the intrinsic character of indistinguishability of identical particles leading to correlations between the events.
II.3 The reality of individual measurement
Ideally one seeks to optimally design experimental setups to mitigate measurement errors and minimize measurement uncertainty. This problem is typically resolved by carrying measurements over a large number of particles with beams of large number of photons to reduce the fluctuations over the measurement. But foundational experiments [14, 15, 16, 19, 17, 18, 20] studying individually one photon or an atom together with separate measurement events each identified as a “click” manifestation in a detector, suggest that a second revolution in quantum mechanics has taken place as one can infer from these experiments that quantum systems can be controlled at the individual level. One easily overlooked matter in this picture is that the so-called “click” signal corresponds in reality to a picture where an amplification of a single-particle signal has been transformed into a macroscopic outcome, as contrary to the latter the former cannot be detected by a human being. Another overlooked point is that, as shown further below, statistics over many events are always used in these “individual” experiments. Claims that each of these events is spatially and timely independent and quantum uncorrelated from each other contradict basic quantum features like uncertainty over position and momentum as well as over time and energy, and the indistinguishability of identical particles. Such claims need to be rigorously examined.
III Continuous measurement versus single-event detection
III.1 Detectors as quantum non-Markovian noise removers
A detector is designed to operate in such a way that the noise of any signal it receives can be reduced or, ideally, suppressed; but the working principle of a detector must also account for the many-body features of an incident beam of photons or atoms. Since a detector is a macroscopic object made of a large number of atoms, the accuracy of any modelling and the validity of the assumptions made should be rigorously assessed. Even though the measurement probability postulate is used in practice in most cases for large numbers of events with the benefit of avoiding a detector modelling, it raises questions about its validity for individual particle detection. A detector does not measure a probability of occurrence; it measures if an event has occurred or not with an uncertainty related to its intrinsic “imperfections”. In particular, for a perfect avalanche diode, realism imposes that any detected quantum photon state does not contain the vacuum state in its superposition. This excludes de facto a photo-detector which transforms by amplification a quantum superposition of a vacuum state and a one-particle state into a cat state superposition of a vacuum state and a many-particle state. If such a process is possible in quantum mechanics through a suitable unitary transformation, it amplifies uncertainty as well. This apparent paradox is overcome by taking into account the spatial extension of a multi-particle and multi-mode wave packet entering into the detector. The detector becomes sensitive once a sufficient amount of a continuous quantum signal has interacted with it to ensure an uncertainty sufficiently small for an effective recording. Therefore, its working principle is non-Markovian not only in the classical sense but also in the quantum sense. Even if the accuracy can be made as high as possible, an unavoidable uncertainty on the particle number or over its energy is always present since a detector is a localized open system. This results in a quantum superposition over particle number and/or energy states.
The detectors models we propose in the following sections, are quite basic to remain tractable but they clearly show how they differ from single-pulse detection modeling.
III.2 Thermodynamics interpretation of the source
Since any particle beam source is always multi-mode and multi-particle, it is crucial to use a quantum field approach for a rigorous analysis of the detection process. Quite generally, the full quantum characterization of a source is impossible since it should be described within the framework of a Fock space by a global many-particle wave function $|\Psi\rangle$. For example, a source of gamma-ray is not point-like but a sample of matter made of a large number of radionuclides emitting gamma photons from wherever they are within the extent of the sample. If a spherical wave from the source can be expected on a statistical average, imperfections due to the finite spatial extension of the source generate a complex superposition of photon states that reveals by interference a localized surge of one or more particles that a detector can subsequently detect unambiguously (like in a bubble chamber or on a photographic plate for instance). Note that for high-energy particles, the uncertainty over the position is in principle very small and makes possible to treat the trajectory classically.
Any particle is described in terms of its quantum field, which for the photon case, around a reference frequency $\omega_{0}$, corresponds to the creation-annihilation operators in the Heisenberg picture $\hat{a}^{\dagger}_{\epsilon}(x,t)$ and $\hat{a}_{\epsilon}(x,t)$ [21], restricting ourselves to the $1+1$ dimensional case of position $x$ and time $t$ for each horizontal or vertical polarisation $\epsilon=H,V$. For a free photon field (we can generalize our reasoning for the more complicated case of a massive particle and/or fermion), they obey the commutation relation on the light cone: $[\hat{a}_{\epsilon}(x,t),\hat{a}^{\dagger}_{\epsilon}(x^{\prime},t^{\prime})]=\delta(x-ct-(x^{\prime}-ct^{\prime}))$. The first-order and second-order correlation functions defined here as:
$g^{(1)}_{\epsilon}(x,t,x^{\prime},t^{\prime})$ $=$ $\langle\Psi|\hat{a}^{\dagger}_{\epsilon}(x,t)\hat{a}_{\epsilon}(x^{\prime},t^{\prime})|\Psi\rangle$
$g^{(2)}_{\epsilon}(x,t,x^{\prime},t^{\prime})$ $=$ $\langle\Psi|\hat{a}^{\dagger}_{\epsilon}(x,t)\hat{a}^{\dagger}_{\epsilon}(x^{\prime},t^{\prime})\hat{a}_{\epsilon}(x,t)\hat{a}_{\epsilon}(x^{\prime},t^{\prime})|\Psi\rangle$
are the source characteristics that are usually known. These correlations can be determined experimentally from statistics over a large particle number. Ideally, a full description should involve higher-order correlations or the photon number per each mode. But even this information would be of no avail to address individually each particle as a consequence of the non-cloning theorem.
In the following, we use the simplification $\hat{a}_{\epsilon}(x,t)=\hat{a}_{\epsilon}(s)$ where $s=x-ct$ is the proper distance, and to proceed, we restrict the spacetime over the pulse interval $L=cT$ with $T$ being the time interval of the localized wave, and $s\in[0,L]$. We then define the global operator $\hat{a}$:
[…]
where we introduce the distribution function $\psi_{\epsilon}(x-ct)=\exp[i\omega_{0}(x/c-t)]f_{\epsilon}(x-ct)$, the quantity $f_{\epsilon}(s)$ being the localized envelope, which satisfies the normalization condition:
[…]
We also define the Fourier transforms:
[…]
$\hat{a}_{\epsilon}(s)$ $=$ $\frac{1}{\sqrt{L}}\sum_{k=-\infty}^{\infty}e^{i(\omega_{0}/c+2\pi k/L)s}\hat{a}_{k,\epsilon}$ (4)
[…]
$f_{\epsilon}(s)$ $=$ $\frac{1}{L}\sum_{k=-\infty}^{\infty}e^{i2\pi ks/L}f_{k,\epsilon}$ (6)
We now focus on the modelling of the radiation considering the following three typical states with very small relative particle number fluctuations:
  1. 1.
    a Fock state with $N$ particles $|N\rangle=(\hat{a}^{\dagger})^{N}|0\rangle/\sqrt{N!}$;
  2. 2.
    a coherent state $|\alpha\rangle=\exp(\alpha\hat{a}^{\dagger}-\alpha^{*}\hat{a})|0\rangle$ with $N=|\alpha|^{2}$;
  3. 3.
    a thermal state with a density matrix
    $\hat{\rho}=\frac{1}{Z}\exp\left[-\beta\sum_{\epsilon=H,V}\sum_{k}E_{k}\hat{a}^{\dagger}_{k,\epsilon}\hat{a}_{k,\epsilon}\right]$ (7)
    where we define $\beta$ as the inverse of an effective temperature, and where $Z$ is the normalisation coefficient. For this last case, we can identify the Bose-Einstein distribution and relate the effective energy $E_{k,\epsilon}$ to the average total particle number $N=\sum_{\epsilon}\langle\hat{N}_{\epsilon}\rangle$. Using the Fourier transform, we find in the weak occupation limit:
    $\langle\hat{a}^{\dagger}_{k,\epsilon}\hat{a}_{k,\epsilon}\rangle\!\!$ $=$ $\!\!N|f_{k,\epsilon}|^{2}\!\!=\!\frac{1}{\exp(\beta E_{k,\epsilon})-1}\!\simeq\!\exp(-\beta E_{k,\epsilon})$ (8)
    $\hat{N}_{\epsilon}\!\!$ $=$ $\!\!\sum_{k=-\infty}^{\infty}\,\hat{a}^{\dagger}_{k,\epsilon}\hat{a}_{k,\epsilon}$ (9)
    $\langle\hat{N}_{\epsilon}\rangle\!\!\!$ $=$ $\!\!\!\!\!\sum_{k=-\infty}^{\infty}\!\!\frac{1}{\exp(\beta E_{k,\epsilon})-1}\!\simeq\!\!\!\sum_{k=-\infty}^{\infty}\!\!\exp(-\beta E_{k,\epsilon})$ (10)
    Quite generally, even using a thermal model where fluctuations are larger than the mean particle number, a deterministic detection is possible with some certainty provided the time window for detection is large enough. The uncertainty is determined from the fluctuations:
    $\langle\delta^{2}\hat{N}_{\epsilon}\rangle=\sum_{k=-\infty}^{\infty}\frac{\exp(\beta E_{k,\epsilon})}{[\exp(\beta E_{k,\epsilon})-1]^{2}}\simeq\langle\hat{N}_{\epsilon}\rangle$ (11)
    This allows to conclude here that the relative fluctuations scale like $\sqrt{\langle\delta^{2}\hat{N}_{\epsilon}\rangle}/\langle\hat{N}_{\epsilon}\rangle\sim 1/\sqrt{N}\sim 1/\sqrt{L}$ and therefore always vanish in the thermodynamic limit. This is the consequence of an average over many modes (like in thermodynamics). The exception is the case of a mono-mode light corresponding to the Bose-Einstein condensation of light with large fluctuations: $\sqrt{\langle\delta^{2}\hat{N}_{\epsilon}\rangle}/\langle\hat{N}_{\epsilon}\rangle\sim 1/N\sim 1/L$, which is not considered here [22].
III.3 Scattering approach
Quite generally, a photon detector can be modelled using scattering theory by analyzing the transition rate of excited electrons using the Fermi golden rule. This rule, which originates from the scattering theory using the S–matrix formalism, imposes an energy conservation between the initial and final states. Although these states look like non-localized states in space such as a pure monochromatic wave, the $i0^{+}$ prescription appearing in the Lippman-Schwinger equations guarantees an adiabatic switching on of a scattering process and imposes that any signal which is interacting with a detector should be causal, localized in time and in space, without any specific characterization of the extension of the wave function. In fact, a localized wave function always has an uncertainty in energy $\Delta E$, which implies to write it as a superposition of delocalized monochromatic input states, which are subsequently transformed into a localized superposition of delocalized output states under the action of S–matrix. As a result of this localisation, the uncertainty principle fixes a lower bound for the time of scattering given by the pulse time $t_{d}\sim\hbar/\Delta E$, which differs from the collision time related to the range of interaction.
Note that we should be cautious to not use the phraseology of transition probabilities as defined for the formal scattering approach [23], as it is valid only once statistics over many events are considered. In contrast for a unique event, the mere assumption of a superposition over many eigenenergy states reestablishes the reality of a time evolution of a localized wave packet for which these rigorous scattering concepts become useless. With this approach, there is a misconception that in a detection of a set of many successive $N$ fainted photon pulses (as for, e.g., a coherent state of very small amplitude, i.e., $|\alpha|^{2}\ll 1$), well separated spatially, they can be assumed independent. Usually, Born’s rule is used to assign a probability of detection for each separated pulse. But since an interaction time can be very long in a scattering process, a photon detection is the result of the accumulation of these many coherent pulses in these detectors as the total intensity $N|\alpha|^{2}$ contains a sufficiently large number of photons. Therefore, the detection process is non-Markovian on the quantum level i.e., a detector cannot reset a virtual state between two successive pulses (unless it is done naturally after a fainted pulse has undergone many reflections) and the single photon detection is a non-instantaneous process resulting from an accumulation of many fainted pulses and not from a probability of occurrence.
III.4 Determinism in photo-detector models
Quite generally, the dependence of the detection efficiency on the (photon or atom) beam intensity is hard to establish especially at the typical low counting rate of 100 photons per second. This raises questions as to the right approach for the calibration of a photo-detector for a single-photon detection. Generally, the black body radiation serves as an intensity reference for the calibration of a photo-detector, which is assumed to have the same sensitivity for any quantum signals such as the thermal radiation or the fainted coherent pulse containing much less than one photon on average. But, the assessment of this assumption requires a microscopic modelling of the detector establishing the transformation process of a single particle signal into a macroscopic current or photo current [21].
In the scattering interpretation, there are two cases:
  1. 1.
    The continuous measurement: We assume a large length $L$ signal with a large average $\langle\hat{N}\rangle\gg 1$ where a detector is in a regime of uncertainty lower than the uncertainty associated to the signal’s noise. The modelling is easy since it simply requires to transform a radiation into a continuous electrical current that is subsequently amplified. The scattering formalism is well suited for a continuous process even for an extremely weak signal as a result of no upper limit for the interaction time [21].
  2. 2.
    Geiger counting: In high energy regimes, here $\geq$ 1 keV, the uncertainty over the energy and hence over the momentum is generally large enough for a particle to be described classically along its trajectory, which can be followed. For instance, an energetic electron or a photon can, along its path in a material medium, excite several thousands of electrons, which can be subsequently analyzed statistically [24]. However, at low energy, here $\sim 1$ eV, it must be ensured that the detector “click” corresponds to one photon only or a defined distribution for the wave function over a window time interval $T$. The discrete measurement is carried out by avalanche photo-diodes. These detectors are difficult to model at the quantum level due to their intrinsic instability. At least, one can assume that the detection is triggered once the certainty for the quantum state of not being the vacuum state, is ensured during the time $T$.
    To give some ideas, we propose three models with different assumptions, showing their non-markovian characteristics:
    Model 1: Considering a coherent input radiation $|\alpha\rangle$ with an average photon number $\langle\hat{N}_{\epsilon}\rangle\sim 1$ and polarisation $\epsilon$, the quantum superposition allows the possibility of including the zero-photon state. Adopting a pragmatic view, we impose that the quantum uncertainty over the absence of photons, measured by $|\langle 0|\alpha\rangle|^{2}=\exp(-\langle\hat{N}_{\epsilon}\rangle)$, is lower than the detector noise or equivalently the inefficiency $1-\eta$ where $\eta$ is the quantum efficiency. As a result, we obtain the working condition $\exp(-\langle\hat{N}_{\epsilon}\rangle)<1-\eta$. For $\langle\hat{N}_{\epsilon}\rangle\sim 2$, the efficiency can reach $85\%$. Such model of a detector is really rudimentary as it neglects its quantum internal dynamics and the possibility of photon losses.
    Model 2: Consider a radiation that induces a Rabi $\pi$-pulse in the Bloch sphere necessary for the detector to operate and excite exactly one electron for an avalanche in the case of a diode, and for a photo-sensitive reaction in the case of a photographic plate. The Rabi pulse time may be smaller than a $\pi$-pulse leading the detector in a state superposition of ground and excited electronic states after interaction with the pulse. The detector works like a ratchet, meaning that many weak pulses produce many small rotations in the Bloch sphere before the avalanche takes place once a complete rotation has been achieved from the south to the north pole. For an effective rotation angle $\theta$, $\pi/\theta$ pulses are necessary to achieve an avalanche. Such a quantumly non-markovian detector memorizes the rotation angle after each pulse with some decoherence occurring between each pulse, thereby reducing effectively this rotation angle. As a consequence, if the time interval between two consecutive pulses is too long, the pulse may scatter away before the next one occurs preventing further rotation and hence the counting rate. An example of such detection process is illustrated in Fig. 1.
    Model 3: The model 2 has the disadvantage that the final state has a well-defined energy in contrast to a semiconductor avalanche detector involving the transition of many electrons between the valence and the conduction bands in the continuum of the wavevectors space. The Fermi golden rule tells that a set of successive distant fainted pulses is decomposed into its wavenumber components as the respective channels of interaction for an effective excitation to take place. The process is therefore collective as it involves many non-separable pulses.In presence of radiation, the conduction electrons number $N_{c}$ obeys the Einstein equation:
    $\frac{dN_{c}}{dt}=-AN_{c}+Bu(N_{v}-N_{c})$ (12)
    where $A$ and $B$ are the Einstein coefficients for the spontaneous and stimulated emission respectively usually calculated using Fermi’s golden rule and $u$ is the spectral energy density per unit bandwidth given by $u=\sum_{\epsilon}\hbar\omega_{0}\langle\hat{N}_{\epsilon}\rangle t_{d}/(Ll^{2})$ with $t_{d}$ the detection time (related to $L=ct_{d}$ for photons) and $l$ the detector size. In thermodynamic equilibrium, the Bose-Einstein statistics associated to the black body radiation must be recovered, which in turns imposes that $A/B=\hbar\omega^{3}_{0}/\pi^{2}c^{3}$. Using similar thermodynamic arguments, the fluctuations $\delta^{2}N_{c}$ obey a dynamics similar to that of $N_{c}$:
    $\delta^{2}N_{c}\sim N_{c}$ (13)
    For a weak constant photon flux and $N_{c}\ll N_{v}$, the spontaneous and stimulated emission rate happen to be negligible. The solution of Eq. (12) for the relative fluctuations scales like $\sqrt{\delta^{2}N_{c}}/N_{c}\sim 1/\sqrt{Bt\sum_{\epsilon}\langle\hat{N}_{\epsilon}\rangle}$, which goes to zero very quickly for large time $t$ guaranteeing a deterministic detection. We can again fix an avalanche detection threshold for $N_{c}$ on the order of unity which can occur after a very long time $t_{d}$ related to the size of the signal.
    Another limitation is the inter-distance between the energy levels $\sim\hbar^{2}/ml^{2}$ within the bands which provides an upper bound to the pulse duration $t_{d}\sim l^{2}m/\hbar$, which is also the detection time. It renders Fermi’s golden rule for estimating $B$ inadequate since the energy band continuum assumption is not valid and would not make the energy transition effective. This point can be illustrated with a simple example. For a photo-diode of cubic size $l=1$ mm and $m=10^{-30}$ kg, we find the upper bound $t_{d}\ll 10$ ms. For larger semi-conductor detectors, this bound becomes even larger. Beyond this bound, the model 2 applies with a decay rate that becomes significant, hence precluding an efficient single-photon detection. Given this time scale and a radiation wavelength of about $\lambda_{0}=2\pi c/\omega_{0}\sim 1\mu m$, we infer from the condition of low spontaneous emission $N_{v}\gg(N_{c}/u)(A/B)$ with $N_{c}=1$ that the minimal carrier density $n_{v}=N_{v}/l^{3}$ in the valence band necessary for an effective transition is:
[…]
    Typically the carrier density is in the 1018 cm-3 range in semiconductors. The existence of such bounds on time and carrier density hinders the isolation of single events well separated in time and with too weak intensity.
These models show the major limitations on the capabilities of detectors, which make it practically impossible with the current available technology to isolate at low energy one pulse event from another in order to control single photons well separated in time. The individual event approach in quantum mechanics works only in the statistical sense. As a result, the advantage of using an avalanche detector in comparison to a continuous detector easier to model, becomes less obvious. The above modelling changes totally the perception of the apparent paradoxes in quantum mechanics, in particular, Schrödinger’s cat and the Bell experiment.
IV The Schrödinger’s cat paradigm
IV.1 Coherent state experiment
The most common state is a coherent state whose uncertainty in the photon number is already present. This state prevents a genuine superposition of a polarisation state as any prism would factorize the resulting state into two disentangled ones:
$|\alpha\rangle$ $=$ $|\alpha_{H}\rangle_{H}|\alpha_{V}\rangle_{V}$ (15)
[…]
The common confusion leading to the Schrödinger’s cat paradox stems from the use of Born’s rule to project Eq. (15) onto a one-particle entangled state, which, assuming a 50%/50% beam splitter, reads
$|\psi\rangle_{E}$ $=$ $(|0\rangle_{H}|1\rangle_{V}+|1\rangle_{H}|0\rangle_{V})/\sqrt{2}$ (17)
[…]
where ${\cal N}_{n}$ enforces the normalisation. The state $|\psi\rangle_{E}$ is the result of a low particle number approximation $|\alpha|\ll 1$, which as a consequence creates a fainted pulse mostly consisting of the vacuum and a one-particle states (states with larger particle numbers being negligible). Equation (17) is an entangled state contrasting with the former, Eq. (15), which is disentangled in the polarization.
[…]
IV.2 The genuine single photon generation
Using a fainted coherent pulse seems a practical way of producing an isolated photon but justifying the absence of effects of higher photon numbers in the superposition of states, is an issue. By exciting a single “artificial” atom by means of a Rabi $\pi$-pulse from a classical radiation source, a subsequent spontaneous emission will produce a real single-photon pulse with a broadening given by the decay rate. Another way to produce one is to use entangled beams where the measurement of one photon in one beam “guarantees” the presence of one photon in another beam [19]. Owing to the precision of an avalanche diode; however, one may question the suitability of such a post-selection method once it is established that there cannot be separability between the events.
If this process is performed within a cavity, the single photon has a channel for producing a genuine superposition leading to a pure entangled state using a beam splitter: $|\psi\rangle_{1}=(c_{V}|0\rangle_{H}|1\rangle_{V}+c_{H}|1\rangle_{H}|0\rangle_{V})$ with $|c_{V}|^{2}+|c_{H}|^{2}=1$ as shown in Fig. 1. Quantum cloning is in principle possible in a known basis and therefore in the Fock basis. If we suppose that an ideal detector leads to the cat state from amplification: $|\psi\rangle_{n}=(c_{V}|0\rangle_{H}|n\rangle_{V}+c_{H}|n\rangle_{H}|0\rangle_{V})$, then the relative fluctuations are very large: $\sqrt{\langle\delta^{2}N_{\epsilon}\rangle}/{\langle\hat{N}_{\epsilon}\rangle}=1/\sqrt{|c_{\epsilon}|^{2}(1-|c_{\epsilon}|^{2})}$, and do not shrink for large cat sizes. But for a realistic case, such states do not appear to be created easily experimentally, since all fluorescence photons cannot be detected individually once we admit that the detection is not a markovian process.
Detector 1 Detector 2
state outcome state outcome
00 00 $\pi/2$ 00
$\pi/2$ 0 $\pi$ or 00 00 or $1$
$\pi$ or 00 00 or $1$ $23\pi/$ 2 or $\pi/2$ 00
$23\pi/$ 2 or $\pi/2$ 00 00 or $\pi$ $1$ or 00
Figure 1: Example of non-Markovian detection based on model 2: Sequence of four $\pi/2$ pulses entering two avalanche detectors, resulting from four separate single photon signals passing through a 50/50 beam splitter (BS). Determinism imposes the absence of a quantum reset, which leads to a memory effect. Once the detector is in a $\pi$ state after the electronic transition, the avalanche is effective or not, according to the detector random noise state. If the two detectors are synchronized, the outcomes are correlated leading to coincidence measurement outcomes. If one detector state is shifted by $\pi/2$ relatively to the other which is more common on average, the outcomes are anti-correlated with no coincidence (see table) and corresponds to the perfect prediction of one photon per pulse if the avalanche is always effective. More generally, the impossibility of the detector to tell its own state due to its noise and the presence of many electrons (see model 3), instead, prevent any synchronisation causing the randomness of both single and coincidence outcomes such as in Ref. [19]. In contrast, a model detector based on Born’s rule would have produced 0/1 or 1/0 outcomes with probability 1/2.
IV.3 Quantum non-demolition measurement
IV.3.1 Superconducting qubit inside a cavity
Quantum non-demolition (QND) measurement is one of the current techniques for monitoring few-level systems (qubit or cavity photon). In a quantum circuit, a qubit readout is achieved by an interaction with a detuned cavity allowing a quantum non-demolition measurement in the dispersive regime. If a qubit is in an initial state $|\psi\rangle_{1}=c_{0}|0\rangle+c_{1}|1\rangle_{V}$, an external input radiation allows, after interaction, the transmission of an ideal coherent state $|\alpha\rangle$ inside a cavity mode for the excited state, the final “cat-like” state $|\psi\rangle_{1}=c_{0}|0\rangle|0\rangle+c_{1}|1\rangle|\alpha\rangle$ [25, 26]. The resulting output signal has two possible unambiguous intensity outcomes 00 or $I=|\alpha|^{2}$ in absence of a superposition. Considering $N$ measurements without a detector reset in between two events, the average total intensity is $\overline{I}=N|c_{1}|^{2}I$ and the fluctuations $\overline{\delta^{2}I}=N[|c_{1}|^{2}(1-|c_{1}|^{2})]I^{2}$. Clearly, the relative fluctuations shrink to zero for large $N$. In this sense, $|c_{0}|^{2}$ and $|c_{1}|^{2}$ are the probability amplitude that is measurable after a counting over many events.
IV.3.2 Rydberg atom passing through a cavity
In Ref. [17], the photon distribution in the state $\sum_{n=1}^{\infty}c_{n}|n\rangle$ is probed in a cavity with a beam of atoms in a Rydberg state, each of which being subsequently detected in one of the two possible spin states, |+ket|$\rangle$| + ⟩ and |ket|$\rangle$| - ⟩, ground and excited, respectively. For each atom individually, a QND operation is carried out on an initially factorized state $|\psi\rangle=(|+\rangle+|-\rangle)\sum_{n=1}^{\infty}c_{n}|n\rangle$ transformed into $\sum_{n=1}^{\infty}[(1+e^{i\theta n)})|+\rangle+(1-e^{i\theta n})|-\rangle]c_{n}|n\rangle$ where $\theta$ is the induced spin rotation angle after the interaction. Starting from a state superposition, successive interrogations induced by successive measurements favor one particular photon number state. Although this experiment suggests that the outcome of the quantum signal originates from the random results of measurements and would generate an effective wavepacket collapse [17], the theory behind this interpretation is not based on a “many-mode” quantum field theory. Again, the assumption over the independence of atoms interacting with the cavity, as they are well spatially separated from each other, cannot work when they are in a superposition of many-body states and may well change the prediction on the cat states.
Since atoms are indistinguishable, we need to work in the many-body representation of a quantum atom field in the two possible Rydberg states. The use of the measurement postulate in a two-state basis for a single atom contrasts with a macroscopic population measurement in two states with the aim of lowering the uncertainty by increasing the atom number. For instance, two atoms evolving into an equal superposition of two Rydberg states can be seen from the detector as one atom in state $g$ and another in state $e$ as they are not separable in time. Therefore, the concept of wave packet collapse is based on the validity of the Born rules to establish the intrinsic randomness of quantum mechanics. A more elaborated model that includes the indistinguishability between atoms would restore the determinism instead and may alter the conclusion established in [17].
V Infeasibility of event separation in the Bell experiments
V.1 General considerations
We neither claim that quantum correlations or entanglement at long distance are not possible, nor that Bell states do not exist. The singlet or triplet states in the Helium atoms are examples of such entangled electron spin pairs. However, performing any Bell-like experiment in a framework based on a many-mode description is questionable once it is established that the collected data correspond to events that are not well separated because of their space and time correlations. Hence, there cannot be independent measurements in the context of a Bell experiment. Such a type of experiment would certainly be feasible if there would be a means to distinguish identical particles, which would substantiate the assumption that two consecutive measurement events can be well separated. But given the experimental reality of non-discrimination between the detector and the setup, the current interpretation of Bell experiments does not involve the physical description of the detector and the source. Without the Born’s postulate, the quantum description of the Bell experiment needs be reformulated. Particularly, the impossibility of a perfect interpretation of a noisy measurement prevents its feasibility, and hence any of two non-local correlated observed outcomes (spin up, spin down) is necessarily the result of many non separable events. Under these conditions, locality, causality, and completeness are restored and there is no need of additional, hidden variables as shown below.
V.2 Influence of the source thermal noise on photon pair production
A set of pair states produced in an experiment is quite different from what John Bell assumed in his work [27]. Using a continuous variable formalism associated to photons, we can factorize the polarized state symmetrically if the EPR source is not noisy. But in this case, all outcomes in both polarisation directions would occur simultaneously contrary to observations where only one direction is observed. A possible explanation for such a discrimination is the inherent thermal noise of the source, like a beam of calcium atoms [14, 15, 16] or a nonlinear crystal [18] that generates additional classical photon flux fluctuations blurring the intrinsic quantum ones, creating a set of random pulse events.
In a particle picture, an itinerant photon shows a Brownian-like dynamics as it interacts with atoms whose motion relative to the detector produces a Doppler shift with a relative frequency change $\sim\sqrt{1/\beta m}/c\sim 10^{-6}$ where $m$ is the mass of the atoms of the source, $c$ the speed of light, and $\beta$ the inverse thermal energy reflecting the temperature statistically determined from the average over the speeds of the atoms in motion.
In addition to the operator $\hat{a}_{k,\epsilon}$, we introduce the pair operator $\hat{b}_{-k,\epsilon}$ both with their specific modes $k,\epsilon$ in one dimension. In the interaction representation, the Hamiltonian describing the source of size $L_{s}$ has the form [28, 29]:
$\hat{H}_{I}(t)$ $=$ $\sum_{k,\epsilon,\epsilon^{\prime}}D_{k,\epsilon,\epsilon^{\prime}}(t)\left(\hat{a}^{\dagger}_{k,\epsilon}\hat{a}_{k,\epsilon^{\prime}}\!+\!\hat{b}^{\dagger}_{k,\epsilon}\hat{b}_{k,\epsilon^{\prime}}\right)$ (19)
$+$ $K\sum_{k,\epsilon}\left(\hat{a}_{k,\epsilon}\hat{b}_{-k,\epsilon}+\hat{a}^{\dagger}_{k,\epsilon}\hat{b}^{\dagger}_{-k,\epsilon}\right)$
where $D_{k,\epsilon,\epsilon^{\prime}}(t)=\Delta_{k}+\eta_{k,\epsilon,\epsilon^{\prime}}(t)$ is decomposed into the detuning term $\Delta_{k}=\omega_{k}-\omega_{p}/2$, which we assume constant in a selected region around the wave number $k_{0}$, and the noise term $\eta_{k,\epsilon,\epsilon^{\prime}}(t)$ due the random Doppler frequency variation. The coupling term $K$ generating the photon pairs, results from the input of two lasers exciting the calcium atoms, or from the non linear crystal with one pump laser of frequency $\omega_{p}$ and amplitude of the pump field $\mathbf{E}$ with an expression scaling like $K\propto|\mathbf{E}|$. It may be time-dependent due to the internal fluctuations of the laser intensity. Under ideal circumstances, the time dependence can be neglected and $\eta_{k,\epsilon,\epsilon^{\prime}}(t)=0$ for a symmetric system under polarization, which means that pairs are produced in all polarisation with equal weight. The crucial point here is that the single value decomposition $D_{k_{0},\epsilon,\epsilon^{\prime}}(t)=\sum_{j=1,2}A^{*}_{\epsilon,j}(t)d_{j}(t)A_{\epsilon^{\prime},j}(t)$ (with $\sum_{\epsilon}|A_{\epsilon,j}|^{2}=1$) provides a history and assigns a preferential polarization when the eigenvalues are not equal $d_{1}(t)\not=d_{2}(t)$.
We can consider it as a stochastic variable whose relative average variation is given by the Doppler shift of a moving atom within the nonlinear crystal or the velocity distribution of calcium atoms. This results in a variance given by $\overline{\eta^{2}_{k,\epsilon,\epsilon^{\prime}}}(t)=\delta_{\epsilon,\epsilon^{\prime}}(\hbar\omega_{k})^{2}/\beta mc^{2}$ Comparing with the detuning and the nonlinear coefficient, these stochastic terms have an effect $(\omega_{k_{0}}/\Delta_{k_{0}})\sqrt{1/\beta mc^{2}}$, which can be of the order of unity or greater. As a consequence, such important contributions lead to potentially strong local variations in the intensity for creating histories of polarization mimicking hidden variables.
A basic formulation of the Bell experiment relies on the multi-mode squeezed state:
$|\Psi(t)\rangle\!=\!\exp\!\!\left[\!\sum_{k,\epsilon,\epsilon^{\prime}=H,V}\!\!\!f_{\epsilon,\epsilon^{\prime}}(t)\hat{a}^{\dagger}_{k,\epsilon}\hat{b}^{\dagger}_{-k,\epsilon^{\prime}}\!-\!f^{*}_{\epsilon,\epsilon^{\prime}}(t)\hat{a}_{k,\epsilon}\hat{b}_{-k,\epsilon^{\prime}}\!\right]\!\!|0\rangle$ (20)
In a steady regime where $D_{k_{0},\epsilon,\epsilon^{\prime}}(t)$ slowly varies in time, we can use the results of Refs. [28, 29] and, based on the Hamiltonian (19), find the connection:
$f_{\epsilon,\epsilon^{\prime}}(t)=\sum_{j=1,2}A^{*}_{\epsilon,j}(t)\lambda_{j}(t)A_{\epsilon^{\prime},j}(t)$ (21)
where $\Omega_{j}(t)=\sqrt{d_{j}^{2}(t)-K^{2}}$ and with the eigenvalues:
$\lambda_{j}(t)\!=\!\frac{K\sin(\Omega_{j}(t)L_{s}/c)}{\left|\cos(\Omega_{j}(t)L_{s}/c)\Omega_{j}(t)\!-\!id_{j}(t)\sin(\Omega_{j}(t)L_{s}/c)\right|}$ (22)
When we trace out over the Fock space associated to the second beam, we recover the form of expression Eq. (7) for the thermal density matrix:
$\hat{\rho}(t)$ $=$ $\frac{1}{Z}\exp\left[\sum_{k,\epsilon,\epsilon^{\prime}=H,V}G_{\epsilon,\epsilon^{\prime}}\hat{a}^{\dagger}_{k,\epsilon}\hat{a}_{k,\epsilon}\right]$ (23)
$G_{\epsilon,\epsilon^{\prime}}$ $=$ $\sum_{j=1,2}A_{\epsilon,j}(t)\ln|\tanh^{2}(\lambda_{j}(t))|A^{*}_{\epsilon^{\prime},j}(t)$ (24)
which guarantees a well-defined average photon number even for a fainted signal, provided many modes $k$ are involved which is the case for a pulse.
Assuming instead that pulses with photon numbers greater than unity are negligible with only one wavevector mode, we could employ the naive approximation of projecting the state onto the particle number pair of Eq. (20) to recover the symmetric Bell entangled state: $|\psi\rangle=(\hat{a}^{\dagger}_{k,H}\hat{b}^{\dagger}_{-k,H}+\hat{a}^{\dagger}_{k,V}\hat{b}^{\dagger}_{-k,V})|0\rangle/\sqrt{2}$. The use of the Born’s probability postulate allows the possibility of different polarisation outcomes.
In comparison using a symmetric state ($\lambda_{1}=\lambda_{2}$) in Eq. (20), it is not possible to discriminate deterministically between two states of polarisation which appear now to be factorized and therefore uncorrelated. With our measurement definition where an avalanche diode is triggered once the number of particles exceeds unity, identical outcomes occur for each beam.
V.3 The more likely realistic case in experiment
Once the probability postulate is dismissed, we can establish that the allegedly “hidden variables” originate from the noise of the sources. These thermal asymmetric fluctuations in the source lead to an asymmetry of polarisation in the wave function favoring one polarisation state over another. For a weak non-linearity $K\ll\Delta_{k_{0}}$, we estimate the relative asymmetry as
$\frac{|\lambda_{1}-\lambda_{2}|}{|\lambda_{1}+\lambda_{2}|}\sim\sqrt{\frac{1}{\beta mc^{2}}}\frac{\omega_{k_{0}}}{\Delta_{k_{0}}}$ (25)
For a fainted pulse, if the total intensity $I\sim\sum_{j}\lambda_{j}^{2}$, then by setting the detector threshold to one half of the total intensity $I_{th}=I/2$, one obtains a majority rule for knowing what polarisation is measured. According to the noise state of the detector, this threshold may also vary, thus leading to an absence of detection or a double counting as an unavoidable error. Under these conditions, the current experiments on Bell inequality may be well interpreted deterministically with an adequate quantum description of the source and the detector.
VI Conclusion
In this work, we critically discussed the Born rule and what its application implies for experimental data interpretation, and consequently for the basic aspects of quantum physics. We argued that once the Born rule as a postulate is dismissed, quantum mechanics can no longer be considered as a probabilistic theory. In fact, this rule is not necessary if a quantum model including both the object of study and the measurement device it interacts with during measurement, based on a scattering approach, is used for data analysis and interpretation. This obviously adds a layer of complexity to the work but with the aim of restoring realism since the existing physical properties of the studied quantum system should not be defined in terms of a probability of occurrence but in terms of macroscopic deterministic quantities. As a consequence, any experiment like the Bell experiment, loses its relevance as a test of the completeness of quantum mechanics, if the assumption on separable single events cannot be fulfilled. Our ignorance of the exact details of the thermal noise in the input state and in the detector state should not imply that quantum mechanics is necessarily probabilistic, but it rather shows the use of probability as a practical convenience. Hence, while Born’s rule may be only applicable as a practical tool for experiments involving a large number of particles leading to a statistical and coarse-grained interpretation, it is unsuitable for experiments involving a few qubits or a few cavity photons to be described as individual entities.
We also discussed with basic illustrative examples the use of the second quantization multi-particle formalism to avoid the pitfalls of the wavepacket collapse in the Schrodinger’s cat paradox. While the use of statistics is inherent to quantum physics experiments, this does not necessary entail that probability is a basic feature of quantum mechanics. It is inherent because the intrinsic uncertainty associated to common observables including, e.g., position, momentum, or particle number, requires a large number of data to be properly defined as a quasi-noiseless macroscopic measurement, which is essential for a conclusive interpretation. This amounts to adopting a thermodynamics formulation of the problem, where only macroscopic quantities are measurable since any single-event signal is too noisy to ensure a valuable characterization. For example, the observation of the electronic cloud in a solid crystal under a scanning tunnelling microscope is not a reality proposed by quantum mechanics but a representation of the total electronic density distribution resulting from of a macroscopic measurement current.
From our proposed many-body modeling including the thermal noise of the source and of the detector, we conclude that the experiments in Refs. [14, 15, 16, 20, 18] are not proven either to correspond to the idealistic Bell’s proposal. Any genuine modeling of a quantum correlation experiment, with the hope of more deterministic predictions, should ideally include: 1) The full formalism of quantum electrodynamics with a many-body formalism highlighting indistinguishability; 2) A correct modelling of the source (e.g., a calcium beam with the two lasers); 3) An accurate quantum model of the detector which operates in a finite time and whose finite size imposes a lower bound over the counting rate preventing genuinely distinguishable (time-separated) single-event detection. A procedure along these steps results in a deterministic quantum description including locality and causality.
More complex experiments that necessitate modeling beyond the traditional scattering approach using the S matrix, such as, e.g., the avalanche effect, need also new theoretical methods to address properly the exact quantum dynamics involved. For instance, the formalism of transmission within wave-guides usually derived from classical electromagnetism has been extended using quantization [21]. These developments raise the fundamental questions on the possibility of creating a superposition of macroscopic “cat” states leading to large quantum fluctuations and therefore not measurable. In this work, we have shown the difficulty to produce such a “multiverse” state given the presently available technology; but a more accurate proof is required to establish this observation on a more formal basis. The non existence of such states amounts to conjecturing that any macroscopic observable has its set of possible eigenvalues restricted to regions within its fluctuations, discarding any extreme value far from its average. Such a statement should ideally be proven starting from a quantum field model.
As a final remark regarding the discrimination in the theoretical description between the object of study and the measurement apparatus, it is worth recalling that when electric circuits were discovered at the time of Alessandro Volta in 19th century, both the circuit (e.g. voltaic pile + metal wires) and the measurement device (e.g. compass or galvanometer) had eventually required the electromagnetic theory for a full understanding of an experiment.
Acknowledgements.
PN is grateful to Mark Everitt, Gilbert Moutaka for fruitful discussions enabling a deeper contextualisation of the problem.